OpenStudy (anonymous):

Evaluate the given expressions. All numbers are approximate. 1328^2-(-7.023)^4/1.037(-2.01) how do i do this? my teachers book got -2256 over -2.08437 and the answer was 1082.341427 which rounds off to 1080. how did she get the top number to be -2256?

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

\[1328^2 - \frac{ -7.023^4 }{ 1.037(-2.01) }\] is this it?

OpenStudy (anonymous):

or everything in numerator

OpenStudy (anonymous):

it's likethe way you have it but she has it like 1328^2-(-7.023)^4/1.037(-2.01)

OpenStudy (anonymous):

i have hw due tomorrow but can't seem to remember this example to do for my hwproblem that is similar

OpenStudy (anonymous):

cause evenwhe n i round i still don't get what she got

OpenStudy (anonymous):

\[\frac{ 2.08437 }{ 2.08437 } \times 1328^2 - \frac{ 7.023^4 }{ 2.08437 } = \frac{ 3,673,528.87 }{ 2.08437 }\] this is what i get. a decimal must be out of place

OpenStudy (anonymous):

yeah that's what i got at home but looking at what we did inclass shows different. but thanks for your help

OpenStudy (anonymous):

i see. it's 1.328^2 answer is: \[\frac{ -2429.035905 }{ 2.08437 }\]

OpenStudy (anonymous):

maybe your teacher rounded up a lot when doing it step by step. im doing it all one step with a calculator. pretty confident about answer. problem is in the problem itself

OpenStudy (anonymous):

ohi have onemore to check it's a bit easier i just want to make sure i solved it right?

OpenStudy (anonymous):

ok np

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay here's my problem...

OpenStudy (anonymous):

x-y^-1/x^-1-y for my final answer i got xy-x-y^2/y is this correct?

OpenStudy (anonymous):

can you please copy/paste it and add brackets?

OpenStudy (anonymous):

can be interpreted so many ways :P

OpenStudy (anonymous):

\[x-y^-1/x^-1-y\]

OpenStudy (anonymous):

your answer is correct

OpenStudy (anonymous):

(xy-x-y^2)/y

OpenStudy (anonymous):

\[\frac{ x-y^-1 }{ x^-1-y}\]

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

oh really. that's not what i read. i read x -[ y^(-1)/x^(-1)] - y

OpenStudy (anonymous):

no bracketes in this just dividing the letter and exponent